OpenStudy (mnr134):

What is the slope of the line with equation y = –x? A. 0 B. 1 C. The slope is undefined. D. –1

rebeccaxhawaii (rebeccaxhawaii):

http://prntscr.com/axov01

rishavraj (rishavraj):

y = mx + c m - slope c - y intercept

rebeccaxhawaii (rebeccaxhawaii):

hint: as you can see the line is decreasing therefore making a negative slope.

OpenStudy (mnr134):

so its 0

rebeccaxhawaii (rebeccaxhawaii):

this would be a zero

rebeccaxhawaii (rebeccaxhawaii):

and also 0 is negative nor positive

OpenStudy (mnr134):

wait so its c.

rebeccaxhawaii (rebeccaxhawaii):

review what i have told you - what is a negative number?

OpenStudy (usukidoll):

. y = -x is a reflection of y = x. since y = x is increasing, y = -x would be decreasing..

OpenStudy (usukidoll):

https://www.montereyinstitute.org/courses/Algebra1/COURSE_TEXT_RESOURCE/U04_L1_T1_text_final_files/image014.gif undefined slope is a vertical line and all points on the line have the same x-coordinate.

OpenStudy (usukidoll):

so undefined slope is out of the picture here... same for the 0 choice because that's just a horizontal line.

OpenStudy (mnr134):

so its d

OpenStudy (usukidoll):

yeah it's a negative slope I'm going to use the slope formula on two points y = -x x = 1,2 y = -1,-2 (1,-1), (2,-2) x1 y1 x2 y2 x2-x1/y2-y1 2-1/-2-(-1) 1/-2+1 1/-1 = -1 which is a negative slope so it's d